The population of country y in 1970 was 60 million. From 1971 to 1975, the total number of births was 25 million and death stood at 3 million. 8 million immigrants and 4 million emigrants were in the reporting period. what is the percentage increase in the population of the country from 1970 to 1975?
1
Expert's answer
2017-11-30T15:40:07-0500
The population of country y in 1970 was 60 million. If from 1971 to 1975, the total number of births was 25 million, death stood at 3 million, 8 million immigrants and 4 million emigrants were in the reporting period, then the every year the population increased by 25 - 3 + 8 - 4 = 26 million. So, there were 60 + 26*5 = 190 million in 1975.
The percentage increase in the population of the country from 1970 to 1975 is (190 - 60)/60*100% = 216.67%.